OpenStudy (anonymous):

Bob has 15,000. He invests it at an interest rate of 4.25 % per years compounded quarterly. a) how much does bob have after one and one half years b)how long will it take his investment to reach 25,000?

OpenStudy (kropot72):

(a) \[A=15000(1+\frac{0.0425}{4})^{(1.5\times4)}=you\ can\ calculate\]

OpenStudy (anonymous):

17,737.59

OpenStudy (anonymous):

Correct? @kropot72

OpenStudy (kropot72):

My calculation gives a different result.

OpenStudy (anonymous):

Oh i had 0.0425/1.5 not 4

OpenStudy (anonymous):

So if I wanted to see how long it will take to reach 25000 I would \[25000=15000(1+\frac{ 0.0425 }{ 4 })^{4\times t}\]

OpenStudy (anonymous):

And for the first you got: 15,982.01?

OpenStudy (kropot72):

Your new result for (a) is correct. Your starting equation to solve (b) is also correct. Good work!

OpenStudy (anonymous):

Lawd have mercy, practice really does make perfect. Thanks dude I might be on tomorrow if i get stuck on anything!

OpenStudy (kropot72):

Lol. I'll help tomorrow if needed.